OpenStudy (seattle12345):

Can some one help with this? Factor: y^3 -4y^2 -25y + 100

OpenStudy (cwrw238):

= y^2(y - 4) - 25(y - 4) = (y^2 - 25)(y - 4) y^2 - 25 is difference to 2 squares so; = (y - 5)( y + 5)(y - 4)

OpenStudy (cwrw238):

- u can often do these by grouping sets of 2 terms

OpenStudy (cwrw238):

do u follow this ok?

OpenStudy (seattle12345):

Yes, I definitely need to practice more though